OpenStudy (anonymous):

Find an equation of a line through (-2,4) and perpendicular to -x+8y=16

OpenStudy (anonymous):

What is the slope of the line -x+8y=16?

OpenStudy (solomonzelman):

okay, at first we want to rearrange your equation into a y-intercept form. \(\large\color{black}{-x+8y=16 }\) \(\large\color{black}{-x+8y\color{red}{+x}=16\color{red}{+x} }\) \(\large\color{black}{8y=16+x }\) then, what do you get after dividing each term (in the equation) by 8?

OpenStudy (anonymous):

y=2=1/8x

OpenStudy (solomonzelman):

yes, you meant, y=2+1/8x , but didn't press shift hard enough the second time:)

OpenStudy (solomonzelman):

So, now we need to find the perpendicular slope. can you do this, or need a hint?

OpenStudy (anonymous):

is it -8

OpenStudy (solomonzelman):

yes, correct.

OpenStudy (solomonzelman):

So now we have, \(\large\color{blue}{ \rm y=-8x+b }\) but we need to find the "b".

OpenStudy (solomonzelman):

you know that (-2,4) is one of the points on your (new) line. (That is a given).

OpenStudy (solomonzelman):

can you solve for b?

OpenStudy (anonymous):

is it -12

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So your equation is now?

OpenStudy (anonymous):

y=-8x-12

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \rm y=-8x+b }\) .... then you found that b=-12.

OpenStudy (solomonzelman):

yes.